Shift report 20011022 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Oct 22 2001 - 16:38:05 EDT

  • Next message: Stephen J. Sanders: "Re: Draft multiplicity paper. BRAHMS 200AGeV."

    ------------------------------------------------------------
    Date: Mon Oct 22 16:38:04 2001, Shift: 08:00-16:00
    Supervisor: JC, with: 
    Log book#10 pages 47 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    1)
    Beam on for almost entire shift.
    
    2)
    !-> Changed D1-D4 polarity to A  before run 5479.
    !-> Changed MRS angle to 40 deg. before run 5481.
    
    DATA runs:
      Run     Trig 6        FS        MRS
      5475     ~120K   4deg, 1/2B    35deg, 4B 
      5479      126K   4deg, 1/2A    35deg, 4B
      5482       15K   4deg, 1/2A    40deg, 4B 
    
    (notice that runs 5482-5483 have MRS at field 4B, 
    3B starts from run 5484)
    
    Pedestals and DVM runs 5476-5478.
    Flemming added capacitor decouplings for ZDC's CFD (run 5482), 
    but this didn't improve the performance, contrary to the 
    expectations, so they were removed.
    
    3)
    DAQ IOServer error ended run 5475. Triggers for this run 
    were not written down to the database (29 seq.).
    
    No detector problems.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Oct 22 2001 - 16:38:44 EDT